The equation of a straight line is given by 2y = 3x - 8 . What is the gradient of the line?
Answer by venugopalramana(3286)

GRADIENT AND SLOPE ARE SAME.SEE THE FOLLOWING EXAMPLE AND DO

-4x-2y=-10 ...the slope(m)and intercept(c) form of a st.line is y=mx+c...so transform the equation into that form
-2y=4x-10
y=(-4/2)x-(10/-2)
y=-2x+5..hence comparing with y=mx+c..we get m=slope=-2 and c=intercept=5